Transfer Pricing, Idle Capacity

Mouton & Perrier, Inc., has a number of divisions that produce liquors, bottled water, and glassware. The Glassware Division manufactures a variety of bottles that can be sold externally (to soft-drink and juice bottlers) or internally to Mouton & Perrier’s Bottled Water Division. Sales and cost data on a case of 24 basic 12-ounce bottles are as follows:

Unit selling price $2.95
Unit variable cost $1.20
Unit product fixed cost* $0.75
Practical capacity in cases 570,000
*$427,500/570,000

During the coming year, the Glassware Division expects to sell 450,000 cases of this bottle. The Bottled Water Division currently plans to buy 106,100 cases on the outside market for $2.95 each. Ellyn Burridge, manager of the Glassware Division, approached Justin Thomas, manager of the Bottled Water Division, and offered to sell the 106,100 cases for $2.90 each. Ellyn explained to Justin that she can avoid selling costs of $0.14 per case by selling internally and that she would split the savings by offering a $0.05 discount on the usual price.

Required:

1. What is the minimum transfer price that the Glassware Division would be willing to accept? Round to the nearest cent.
$ per unit

What is the maximum transfer price that the Bottled Water Division would be willing to pay? Round to the nearest cent.
$ per unit

Should an internal transfer take place?

What would be the benefit (or loss) to the firm as a whole if the internal transfer takes place? When required, round your answer to the nearest dollar.
  $

2. Suppose Justin knows that the Glassware Division has idle capacity. Do you think that he would agree to the transfer price of $2.90?

Suppose he counters with an offer to pay $2.42. If you were Ellyn, would you be interested in this price?

3. Suppose that Mouton & Perrier’s policy is that all internal transfers take place at full manufacturing cost. What would the transfer price be? Round to the nearest cent.
$ per unit

Would the transfer take place?
Answer #1

1.Since there is spare capacity, minimum price is equal to the variable cost incurred

= 1.20-0.14

= $1.06 per unit

2.Maximum Price is equal to the market Price i.e. 2.95 per unit

Yes

Benefit to the firm = (2.95-1.06)*106,100 = $200,529

2. No

Yes, as it is still higher than variable cost

3.Transfer price = 1.20-0.14+0.75

= $1.81 per unit

Yes